# Mixed Warm Dark Matter Constraints using Milky Way Satellite Galaxy Counts

### v2

@@ I Introduction
Numerous studies place strict constraints on WDM as the sole component of dark matter for both the general thermal relic case and specific sterile neutrino models (Zelko et al., 2022; Lovell, 2023). Using MW satellite counts, various works have placed stringent constraints on the mass of thermal relic WDM (Kennedy et al., 2014; Lovell et al., 2016; Enzi et al., 2021a; Newton et al., 2021; Dekker et al., 2022; Newton et al., 2024), the strongest of which yields a 95% Bayesian lower limit of $m_{\rm WDM}\geq$ 6.5 keV (Nadler et al., 2021). Other measurements of the Lyman–$\alpha$ forest (e.g., Iršič et al., 2024), UV luminosity function (e.g., Liu et al., 2024), stellar streams (e.g., Banik et al., 2021) and gravitationally lensed quasars (e.g., Keeley et al., 2024) have also been used to set limits on the thermal relic WDM mass at $m_{\rm WDM}\geq$ 5.7, 3.2, 3.6, and 6.1 keV, respectively (using a variety of significance criteria). Moreover, through a combined analysis of MW satellites counts and strong gravitational lenses, Nadler et al. (2021) produced a stringent joint constraint of $m_{\rm WDM}\geq$ 9.7 keV at a 95% Bayesian lower limit and 7.4 keV at a 20:1 marginal likelihood ratio. While Enzi et al. (2021b) obtained limits of $m_{\rm WDM}\geq$ 6.0 keV at a 95% Bayesian lower limit and 2.6 keV at a 20:1 likelihood ratio by combining the analysis of gravitational lenses, Lyman-$\alpha$ forest, and MW satellite counts.
@@
In this analysis, we use the population counts MW satellite galaxies found in Dark Energy Survey (DES: Abbott et al., 2016) and Pan-STARRS1 (PS1: Chambers et al., 2016) survey to put constraints on MWDM properties. In Section II, we first discuss the Semi-Analytical Sub-Halo Inference ModelIng code (sashimi; Hiroshima et al., 2018; Ando et al., 2019; Dekker et al., 2022) used to obtain the mass distribution of subhalos for different MWDM scenarios and present an analytical approximation for the suppression of the subhalo mass function relative to CDM. Section III describes the galaxy–halo connection model (Nadler et al., 2021) used to predict the number of MW satellites that would be observable in different MWDM scenarios. We compare these predictions to the observed MW satellite counts to obtain constraints on mixed thermal-relic scenarios (Section IV) and mixed sterile neutrinos scenarios (Section V). We discuss our analysis and conclude in Section VI. In our analysis, we adopt the flat $\Lambda$MWDM model as our fiducial cosmology with $H_{0}=70$ km s-1 Mpc-1, $h=0.7$, $\Omega_{\rm m}=0.286$, $\Omega_{\Lambda}=0.714$, $\Omega_{\rm b}=0.047$ , $\sigma_{8}=0.82$, $n_{s}=0.96$.
@@ II Subhalo population of Mixed Thermal-relic WDM
We account for the MWDM scenario by multiplying the distribution of subhalo masses in CDM, known as the subhalo mass function, with a subhalo suppression function $\beta(M,m_{\rm WDM},f_{\rm WDM})$, defined as
@@
where $dN_{\rm sub}/{dM}$ is the subhalo mass function, $M$ is the subhalo mass, $\theta_{\rm wdm}$ is the model-specific properties of the WDM scenario such as its particle mass, and $f_{\rm WDM}$ is the WDM fraction. Throughout this work, we calculate the subhalo suppression function based on the subhalo mass at accretion within a radius where its mean density is 200 times the critical density of the Universe, which we assume is equal to its peak mass.
@@
In order to obtain the suppression function for a large range of WDM scenarios, we use the semi-analytical sashimi code (Hiroshima et al., 2018; Ando et al., 2019; Dekker et al., 2022). sashimi predicts subhalo properties for a given matter power spectrum and host halo mass, and has been tested against numerical simulations in the case of pure CDM and WDM. Semi-analytical models provide a fast and flexible platform for computing subhalo properties, allowing us to make quick estimates of the predictions from different dark matter scenarios. Alternative approaches exploring MWDM scenarios through cosmological N-body simulations are forthcoming and will provide a valuable cross check (An et al., 2024).
@@
For comparison, Lovell et al. (2014) found the following fitting parameters in the case of pure WDM: $\alpha=1.35(m_{\rm WDM}/3\rm{keV})^{-10/3}$, $\gamma=-0.99$, and $\Delta=1$.
Figure 1: The subhalo suppression function, $\beta(M,m_{\rm WDM},f_{\rm WDM})$, for thermal-relic WDM with mass of 2 keV (left), 5 keV (middle), 8 keV (right) as a function of subhalo mass $M$. The colored solid lines show the suppression functions for different fractions, $f_{\rm WDM}$, from sashimi, while the black dashed lines show the best-fit analytical suppression functions (see Eq. 4.)
@@
where $i$ indexes the simulated subhalos, $f_{\rm gal}$ is the galaxy occupation fraction (probability that the subhalo will host a satellite galaxy), $p_{\rm disrupt}$ is the probability that a subhalo would be disrupted due to baryonic effects (Nadler et al., 2018), $p_{\rm detect}$ is probability of the galaxy being detected in DES and PS1 calculated using survey selection functions from Drlica-Wagner et al. (2020), and $\beta$ (Eq. 4) is the subhalo suppression function relative to the simulated subhalo population in CDM. For a given set of galaxy–halo connection parameters ($\theta_{\rm model}=\{\alpha,\sigma_{M},\mathcal{A},n,\sigma_{\log R},\mathcal{M}_{50},\sigma_{\rm gal},\mathcal{B}\}$, see Nadler et al. (2020) for more details), and a specific subhalo suppression function, $\beta$, we can generate a predicted population of satellite galaxies.
@@ IV Limits on Mixed Thermal Relic WDM
To incorporate the luminosity and size information of the satellites, we bin the predicted and observed satellites based on their absolute magnitude and surface brightness. [^2] We first bin the galaxies into 7 equally-spaced bins based on their absolute magnitude from $200\geq M_{V}\geq-$ 20. We further split the galaxies into 4 bins, based on a surface brightness threshold at 28 mag arcsec-2 and whether the galaxies would be found in the DES and PS1, resulting in a total of 28 galaxy bins. We obtain the likelihood that the predicted number of galaxies for each MWDM scenario represent the observed MW satellite population in DES and PS1 by assuming that the predicted number of satellite galaxies follows a Poisson distribution in each bin (Nadler et al., 2020).
Fig. 2 shows the binned predicted luminosity functions of galaxies observed in both the DES and PS1 regions for CDM and MWDM comprising of thermal-relic WDM with $m_{\rm WDM}=2$ keV. For each scenario, we use the Markov Chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013) to sample the posterior distribution of the galaxy–halo connection parameters. The shaded band represents the uncertainty due to different galaxy-halo connection model values and the stochasticity of the model. The uncertainties of the luminosities and number counts of the observed satellites is associated with the size of the luminosity bins used in the analysis and Poisson statistics, respectively. We find that due to the uncertainty in the galaxy-halo connection parameters, the $f_{\rm WDM}=0.2$, $m_{\rm WDM}=2$ keV mixed thermal-relic WDM scenario predicts roughly the same number of MW satellite galaxies as CDM, while the $f_{\rm WDM}=0.6$ scenario starts to show a deviation from CDM. Note that for illustrative purposes, we only bin the galaxies based on their luminosities in Fig. 2.
@@
We use emcee (Foreman-Mackey et al., 2013) to sample the marginalized posterior distribution of the galaxy–halo connection model for different mixed thermal-relic WDM models. In our analysis, we scan over fixed WDM mass and leave the WDM fraction as a free parameter between $0\leq f_{\rm WDM}\leq 1$. For each WDM mass, we run emcee for 20,000 steps with 8 walkers to sample the posterior of nine free parameters in the model, $\theta_{\rm mcmc}=\{\alpha,\sigma_{M},\mathcal{A},n,\sigma_{\log R},\mathcal{M}_{50},\sigma_{\rm gal},\mathcal{B},f_{\rm WDM}\}$. The first eight parameters are the galaxy–halo connection parameters, and we marginalize over them in this analysis. We run the MCMC chains for WDM masses in the range of $1.5\leq m_{\rm WDM}\leq 8.0$ keV at mass intervals of 0.5 keV.
@@
An alternative method commonly quoted in the WDM literature is the 95% credible interval (Viel et al., 2013; Nadler et al., 2021), which is defined as the range of parameter values which enclosed 95% of the integrated marginal posterior. While the credible interval methods includes more information of the shape of the posterior, our posteriors are unbounded in the limit of large $m_{\rm WDM}$, and thus the limits obtained would be sensitive to the boundary of the priors adopted for the WDM parameters which can be arbitrary. [^3] The full posterior distribution can be found https://github.com/chinyitan/mwdm.
Fig. 3 shows the parameter space that is disfavored by the MW satellite galaxy counts. We exclude thermal-relic WDM with mass $m_{\rm WDM}\lesssim 6.6$ keV (5.8 keV) with a posterior ratio of 10:1 (or 20:1) from being all of dark matter. This is similar to the pure WDM constraints found in Nadler et al. (2021), which excluded $m_{\rm WDM}\leq 5.2$ keV at marginalized posterior ratio of 20:1 using a slightly different form for the suppression of the subhalo mass function. As we move towards lower $m_{\rm WDM}$, our allowed WDM fraction, $f_{\rm WDM}$, decreases until our constraints starts to converge at $f_{\rm WDM}\sim 0.45$ (0.50). Moreover, we find that fixing the galaxy–halo connection parameters would result in a constraint on $f_{\rm WDM}$ that linearly decreases with $m_{\rm WDM}$, lowering the minimum allowed $f_{\rm WDM}$ values below $f_{\rm WDM}\sim 0.45$ for $m_{\rm WDM}\lesssim 2$ keV. In Fig. 3, we also compare our constraints with previous constraints, as shown as grey shaded regions from gravitational lenses (Kamada et al., 2016), and cluster number counts (Parimbelli et al., 2021). Additionally, a similar analysis constraining MWDM models by comparing MW satellite observations to cosmological N-body zoom-in simulations yields qualitatively similar results (An et al., 2024), lending strength to both approaches. Moreover, the star represent constraints from the 21-cm absorption feature claimed by EDGES (Bowman et al., 2018), which excludes $f_{\rm WDM}\geq 0.17$ for $m_{\rm WDM}\sim$ 1 keV (Schneider, 2018).
@@ V Limits on Mixed Sterile Neutrino
We use the public sterile-dm code (Venumadhav et al., 2016) to generate the phase-space distributions of $\nu_{s}$ and $\bar{\nu}_{s}$ for a range of lepton asymmetry values that correspond to sterile neutrino abundances, $\Omega_{\nu_{s}}$, and mixing angles, $\sin^{2}(2\theta)$. We obtain the shape of the matter power spectrum by providing the phase-space distributions as an input for the Boltzmann code CLASS (Lesgourgues, 2011). We obtain the subhalo suppression function using sashimi for mixing angles and fractions in the range of ${4\times 10^{-13}}\leq\sin^{2}(2\theta)\leq 10^{-11}$ and $0.1\leq f_{\nu_{s}}\leq 0.6$ with fraction steps of 0.1. The subhalo suppression function can be approximated using Eq. 4, for which we fit the parameters $\alpha$, $\gamma$, and $\Delta$ for fractions in the range of $0.1\leq f_{\nu_{s}}\leq 0.6$.  [^4] The subhalo suppression functions for the mixed sterile neutrino scenerios considered in this analysis can be found https://github.com/chinyitan/mwdm.
@@ VI Discussion and Conclusion
Using the population of MW satellite galaxies detected by DES and PS1, we put new constraints on MWDM scenarios. We use the CDM subhalo population obtained from Mao et al. (2015) and implemented the MWDM scenario by suppressing the subhalo population using an analytical function, $\beta(M,m_{\rm WDM},f_{\rm WDM})$, derived from the semi-analytical subhalo inference model, sashimi (Hiroshima et al., 2018; Ando et al., 2019; Dekker et al., 2022). Using the galaxy–halo connection model from Nadler et al. (2020, 2021), we estimate the predicted number of MW satellites for different MWDM scenarios and calculate the likelihood of observing the known satellites in DES and PS1 given the MWDM model parameters.
We find that MW satellite galaxy counts constrain the fraction of thermal WDM at $m_{\rm WDM}\sim 6.6$ keV for $f_{\rm WDM}=1.0$ at a posterior ratio threshold of 10:1. As we decrease the WDM mass, our allowed WDM fraction decreases to $f_{\rm WDM}\sim 0.45$ for $m_{\rm WDM}\sim 1.5$ keV. These constraints from MW satellite counts extend existing MWDM constraints to much higher masses ($m_{\rm WDM}>1$ keV) compared to previous analysis that used different methods. Due to uncertainties in the galaxy–halo connection parameters and the limited population of observed MW satellites, our analysis cannot differentiate among MWDM scenarios with $f_{\rm WDM}\lesssim 0.45$ leading our constraints to asymptote to this value at low $m_{\rm WDM}$.
@@
In this work, we relied on only two cosmological zoom-in simulations for our MW subhalo population. This limits our ability to investigate uncertainties in the host halo mass (which changes the normalization of the subhalo mass function and thus the total number of subhalos), host environments and formation histories. Recent efforts to simulate a much larger suite of MW-like hosts in CDM (Nadler et al., 2023; Buch et al., 2024) and alternative dark matter scenarios (Nadler et al., 2024a; An et al., 2024) are expected to better constrain these systematic uncertainties. In addition, sashimi could be used to generate subhalo populations for a range of host halo masses; however, this would require modifications to incorporate spatial information of the subhalo population and the effects of the LMC.
Our constraints come from the MW satellite population found in DES and PS1 and the corresponding survey selection functions (Drlica-Wagner et al., 2020). However, the known population of MW satellites has grown significantly over the past few years (e.g., Mau et al., 2020; Cerny et al., 2021, 2023a, 2023b; Smith et al., 2023; Homma et al., 2024; Tan et al., 2024). Furthermore, current and near-future surveys such as the DECam Local Volume Exploration survey (DELVE; Drlica-Wagner et al., 2021, 2022), the Hyper Suprime-Cam Strategic Survey Program (HSC-SSP; Aihara et al., 2018), the Ultraviolet Near Infrared Optical Northern Survey (UNIONS; Ibata et al., 2017), and the Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) are expected significantly increase the known population of MW satellites in the coming years (e.g., Hargis et al., 2014; Jethwa et al., 2018; Newton et al., 2018; Nadler et al., 2020; Manwadkar and Kravtsov, 2022; Nadler et al., 2024b). This growing population of UFDs will improve our understanding of dark matter and allow for more sensitive searches for potential deviation from CDM.

### v3

@@ I Introduction
Numerous studies place strict constraints on WDM as the sole component of dark matter for both the general thermal relic case and specific sterile neutrino models (Zelko et al., 2022; Lovell, 2023). Using MW satellite counts, various works have placed stringent constraints on the mass of thermal relic WDM (Kennedy et al., 2014; Lovell et al., 2016; Enzi et al., 2021a; Newton et al., 2021; Dekker et al., 2022; Newton et al., 2024), the strongest of which yields a 95% Bayesian lower limit of $m_{\rm WDM}\geq$ 6.5 keV (Nadler et al., 2021). Other measurements of the Lyman–$\alpha$ forest (e.g., Iršič et al., 2024), UV luminosity function (e.g., Liu et al., 2024), stellar streams (e.g., Banik et al., 2021) and gravitationally lensed quasars (e.g., Keeley et al., 2024) have also been used to set limits on the thermal relic WDM mass at $m_{\rm WDM}\geq$ 5.7, 3.2, 3.6, and 6.1 keV, respectively (using a variety of significance criteria). Moreover, through a combined analysis of MW satellites counts and flux ratios of quadruply-lensed quasars, Nadler et al. (2021) produced a stringent joint constraint of $m_{\rm WDM}\geq$ 9.7 keV at a 95% Bayesian lower limit and 7.4 keV at a 20:1 marginal likelihood ratio. While Enzi et al. (2021b) obtained limits of $m_{\rm WDM}\geq$ 6.0 keV at a 95% Bayesian lower limit and 2.6 keV at a 20:1 likelihood ratio by combining the analysis of resolved imaging of galaxy–galaxy lenses, Lyman-$\alpha$ forest, and MW satellite counts.
@@
In this analysis, we compare predictions for the luminosity functions of MW satellite galaxies from various MWDM models with the MW satellite galaxies observed in the Dark Energy Survey (DES: Abbott et al., 2016) and the Pan-STARRS1 (PS1: Chambers et al., 2016) survey to set constraints on MWDM properties. In Section II, we first discuss the Semi-Analytical Sub-Halo Inference ModelIng code (sashimi; Hiroshima et al., 2018; Ando et al., 2019; Dekker et al., 2022) used to obtain the mass distribution of subhalos for different MWDM scenarios and present an analytical approximation for the suppression of the subhalo mass function relative to CDM. Section III describes the galaxy–halo connection model (Nadler et al., 2021) used to predict the number of MW satellites that would be observable in different MWDM scenarios. We compare these predictions to the observed MW satellite counts to obtain constraints on mixed thermal-relic scenarios (Section IV) and mixed sterile neutrinos scenarios (Section V). We discuss our analysis and conclude in Section VI. In our analysis, we adopt the flat $\Lambda$MWDM model as our fiducial cosmology with $H_{0}=70$ km s-1 Mpc-1, $h=0.7$, $\Omega_{\rm m}=0.286$, $\Omega_{\Lambda}=0.714$, $\Omega_{\rm b}=0.047$ , $\sigma_{8}=0.82$, $n_{s}=0.96$.
@@ II Subhalo population of Mixed Thermal-relic WDM
We account for the MWDM scenario by multiplying the distribution of subhalo masses in CDM, known as the subhalo mass function, with a subhalo suppression function, $\beta(M,\theta_{\rm WDM},f_{\rm WDM})$, such that
@@
where $dN_{\rm sub}/{dM}$ is the subhalo mass function, $M$ is the subhalo mass, $\theta_{\rm WDM}$ is the model-specific parameters of the WDM scenario such as its particle mass, and $f_{\rm WDM}$ is the WDM fraction. Throughout this work, we calculate the subhalo suppression function based on the subhalo mass at accretion within a radius where its mean density is 200 times the critical density of the Universe, which we assume is equal to its peak mass.
@@
In order to obtain the suppression function for a large range of WDM scenarios, we use the semi-analytical sashimi code (Hiroshima et al., 2018; Ando et al., 2019; Dekker et al., 2022). sashimi predicts subhalo properties for a given matter power spectrum and host halo mass, and has been tested against numerical simulations in the case of pure CDM and WDM. Specifically for this work, we have verified that the CDM subhalo mass function generated by sashimi is consistent with the subhalo populations obtained from the two simulations by Mao et al. (2015). Semi-analytical models provide a fast and flexible platform for computing subhalo properties, allowing us to make quick estimates of the predictions from different dark matter scenarios. Alternative approaches exploring MWDM scenarios through cosmological N-body simulations are forthcoming and will provide a valuable cross check (An et al., 2024).
@@
For comparison, Lovell et al. (2014) found the following fitting parameters in the case of pure WDM: $\alpha\approx 52.6(m_{\rm WDM}/1\rm{keV})^{-3.33}$, $\gamma=-0.99$, and $\Delta=1$, assuming that the half-mode mass is given by $M_{\rm hm}=5\times 10^{8}(m_{\rm WDM}/3\rm{keV})^{-10/3}M_{\odot}$(Nadler et al., 2021).
Figure 1: The subhalo suppression function, $\beta(M,m_{\rm WDM},f_{\rm WDM})$, for thermal-relic WDM with mass of 2 keV (left), 5 keV (middle), 8 keV (right) as a function of subhalo mass $M$. The colored solid lines show the suppression functions for different fractions, $f_{\rm WDM}$, from sashimi, while the black dashed lines show the best-fit analytical suppression functions (see Eq. 4). Also shown is the pure thermal-relic WDM ($f_{\rm WDM}$=1) subhalo suppression function from Lovell et al. (2014)
@@
where $i$ indexes the simulated subhalos, $f_{\rm gal}$ is the galaxy occupation fraction (probability that the subhalo will host a satellite galaxy), $p_{\rm disrupt}$ is the probability that a subhalo would be disrupted due to baryonic effects (Nadler et al., 2018), $p_{\rm detect}$ is probability of the galaxy being detected in DES and PS1 calculated using survey selection functions from Drlica-Wagner et al. (2020), and $\beta$ (Eq. 4) is the subhalo suppression function relative to the simulated subhalo population in CDM. For a given set of galaxy–halo connection model parameters ($\theta_{\rm model}=\{\alpha,\sigma_{M},\mathcal{A},n,\sigma_{\log R},\mathcal{M}_{50},\sigma_{\rm gal},\mathcal{B}\}$, see Nadler et al. (2020) for more details), and a specific subhalo suppression function, $\beta$, we can generate a predicted population of satellite galaxies.
@@ IV Limits on Mixed Thermal Relic WDM
To incorporate the luminosity and size information of the satellites, we bin the predicted and observed satellites based on their absolute magnitude and surface brightness  [^2] We first bin the galaxies into 6 equally-spaced bins based on their absolute magnitude from $200\geq M_{V}\geq-$ 20. We further split the galaxies into 4 bins, based on a surface brightness threshold at 28 mag arcsec-2 and whether the galaxies would be found in the DES and PS1, resulting in a total of 24 galaxy bins.. We obtain the likelihood that the predicted number of galaxies for each MWDM scenario represent the observed MW satellite population in DES and PS1 by assuming that the predicted number of satellite galaxies follows a Poisson distribution in each bin (Nadler et al., 2020).
Fig. 2 shows the binned predicted luminosity functions of galaxies observed in both the DES and PS1 regions for CDM and MWDM comprising of thermal-relic WDM with $m_{\rm WDM}=2$ keV. For each scenario, we use the Markov Chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013) to sample the posterior distribution of the galaxy–halo connection model parameters. The shaded band represents the uncertainty due to different galaxy–halo connection model values and the stochasticity of the model. The uncertainties of the luminosities and number counts of the observed satellites is associated with the size of the luminosity bins used in the analysis and Poisson statistics, respectively. We find that due to the uncertainty in the galaxy–halo connection model parameters, the $f_{\rm WDM}=0.2$, $m_{\rm WDM}=2$ keV mixed thermal-relic WDM scenario predicts roughly the same number of MW satellite galaxies as CDM, while the $f_{\rm WDM}=0.6$ scenario starts to show a deviation from CDM. Note that for illustrative purposes, we only bin the galaxies based on their luminosities in Fig. 2.
@@
We use emcee (Foreman-Mackey et al., 2013) to sample the marginalized posterior distribution of the galaxy–halo connection model for different mixed thermal-relic WDM models. In our analysis, we scan over fixed WDM mass and leave the WDM fraction as a free parameter between $0\leq f_{\rm WDM}\leq 1$. For each WDM mass, we run emcee for 20,000 steps with 8 walkers to sample the posterior of nine free parameters in the model, $\theta_{\rm mcmc}=\{\alpha,\sigma_{M},\mathcal{A},n,\sigma_{\log R},\mathcal{M}_{50},\sigma_{\rm gal},\mathcal{B},f_{\rm WDM}\}$. The first eight parameters are the galaxy–halo connection model parameters, and we marginalize over them in this analysis. We run the MCMC chains for WDM masses in the range of $1.5\leq m_{\rm WDM}\leq 8.0$ keV at mass intervals of 0.5 keV.
@@
An alternative method commonly quoted in the WDM literature is the 95% credible interval (Viel et al., 2013; Nadler et al., 2021), which is defined as the range of parameter values which enclosed 95% of the integrated marginal posterior. While the credible interval methods includes more information of the shape of the posterior, our posteriors are unbounded in the limit of large $m_{\rm WDM}$, and thus the limits obtained would be sensitive to the boundary of the priors adopted for the WDM parameters which can be arbitrary  [^3] The full posterior distribution can be found https://github.com/chinyitan/mwdm..
Fig. 3 shows the parameter space that is disfavored by the MW satellite galaxy counts. We exclude thermal-relic WDM with mass $m_{\rm WDM}\lesssim 6.6$ keV (5.8 keV) with a posterior ratio of 10:1 (or 20:1) from being all of dark matter. This is similar to the pure WDM constraints found in Nadler et al. (2021), which excluded $m_{\rm WDM}\leq 5.2$ keV at marginalized posterior ratio of 20:1 using a slightly different form for the suppression of the subhalo mass function. As we move towards lower $m_{\rm WDM}$, our allowed WDM fraction, $f_{\rm WDM}$, decreases until our constraints starts to converge at $f_{\rm WDM}\sim 0.45$ (0.50).
In Fig. 3, we also compare our constraints with previous constraints, as shown as grey shaded regions from gravitational lenses (Kamada et al., 2016), and cluster number counts (Parimbelli et al., 2021). Additionally, a similar analysis constraining MWDM models by comparing MW satellite observations to cosmological N-body zoom-in simulations yields qualitatively similar results (An et al., 2024), lending strength to both approaches. Moreover, the star represent constraints from the 21-cm absorption feature claimed by EDGES (Bowman et al., 2018), which excludes $f_{\rm WDM}\geq 0.17$ for $m_{\rm WDM}\sim$ 1 keV (Schneider, 2018).
In this analysis, we adopted the same wide priors on the galaxy–halo connection model parameters as Nadler et al. (2020). As a result, in some MWDM scenarios, the suppression in the subhalo mass function can be compensated by adjusting the values of the galaxy–halo connection model parameters. This, in turn, weakens our MWDM constraints. For example, we find that the posterior distribution of the faint-end slope of the subhalo-satellite luminosity function, $\alpha$, in our MWDM scenarios is lower compared to CDM. The faint-end slope determines the luminosity of a satellite galaxy as a function of subhalo peak circular velocity. A lower value of $\alpha$ allows less massive subhaloes to produce brighter satellites, making these galaxies more easily detectable and consequently increasing the predicted number of detected galaxies. We also observe similar effects in the other galaxy–halo connection model parameters but to a much smaller extent. A more extensive discussion of the galaxy–halo connection model can be found in Nadler et al. (2019b) and Nadler et al. (2020)
@@ V Limits on Mixed Sterile Neutrino
We use the public sterile-dm code (Venumadhav et al., 2016) to generate the phase-space distributions of $\nu_{s}$ and $\bar{\nu}_{s}$ for a range of lepton asymmetry values that correspond to sterile neutrino abundances, $\Omega_{\nu_{s}}$, and mixing angles, $\sin^{2}(2\theta)$. We obtain the shape of the matter power spectrum by providing the phase-space distributions as an input for the Boltzmann code CLASS (Lesgourgues, 2011). We obtain the subhalo suppression function, $\beta(M,\sin^{2}(2\theta),f_{\nu_{s}})$, using sashimi for mixing angles and fractions in the range of ${4\times 10^{-13}}\leq\sin^{2}(2\theta)\leq 10^{-11}$ and $0.1\leq f_{\nu_{s}}\leq 0.6$ with fraction steps of 0.1. The subhalo suppression function can be approximated using Eq. 4, for which we fit the parameters $\alpha$, $\gamma$, and $\Delta$ for fractions in the range of $0.1\leq f_{\nu_{s}}\leq 0.6$  [^4] The subhalo suppression functions for the mixed sterile neutrino scenerios considered in this analysis can be found https://github.com/chinyitan/mwdm..
@@ VI Discussion and Conclusion
Using the population of MW satellite galaxies detected by DES and PS1, we put new constraints on MWDM scenarios. We use the CDM subhalo population obtained from Mao et al. (2015) and implemented the MWDM scenario by suppressing the subhalo population using an analytical function, $\beta(M,\theta_{\rm WDM},f_{\rm WDM})$, derived from the semi-analytical subhalo inference model, sashimi (Hiroshima et al., 2018; Ando et al., 2019; Dekker et al., 2022). Using the galaxy–halo connection model from Nadler et al. (2020, 2021), we estimate the predicted number of MW satellites for different MWDM scenarios and calculate the likelihood of observing the known satellites in DES and PS1 given the MWDM model parameters.
We find that MW satellite galaxy counts constrain the fraction of thermal WDM at $m_{\rm WDM}\sim 6.6$ keV for $f_{\rm WDM}=1.0$ at a posterior ratio threshold of 10:1. As we decrease the WDM mass, our allowed WDM fraction decreases to $f_{\rm WDM}\sim 0.45$ for $m_{\rm WDM}\sim 1.5$ keV. These constraints from MW satellite counts extend existing MWDM constraints to much higher masses ($m_{\rm WDM}>1$ keV) compared to previous analysis that used different methods. Due to uncertainties in the galaxy–halo connection model parameters and the limited population of observed MW satellites, we are not able to constrain MWDM fractions below $f_{\rm WDM}\lesssim 0.45$ for any $m_{\rm WDM}$ we consider. However, Nadler et al. (2024a) have demonstrated that with a larger population of fainter satellites, it will be possible to better disentangle the effects of WDM and the uncertainties in the galaxy–halo connection model parameters. This would therefore allow us to simultaneously achieve better constraints on galaxy–halo connection model parameters and to better distinguish between different dark matter models.
@@
In this work, we relied on only two cosmological zoom-in simulations for our MW subhalo population. This limits our ability to investigate uncertainties in the host halo mass (which changes the normalization of the subhalo mass function and thus the total number of subhalos), host environments and formation histories. Recent efforts to simulate a much larger suite of MW-like hosts in CDM (Nadler et al., 2023; Buch et al., 2024) and alternative dark matter scenarios (Nadler et al., 2024b; An et al., 2024) are expected to better constrain these systematic uncertainties. In addition, sashimi could be used to generate subhalo populations for a range of host halo masses; however, this would require modifications to incorporate spatial information of the subhalo population and the effects of the LMC.
Our constraints come from the MW satellite population found in DES and PS1 and the corresponding survey selection functions (Drlica-Wagner et al., 2020). However, the known population of MW satellites has grown significantly over the past few years (e.g., Mau et al., 2020; Cerny et al., 2021, 2023a, 2023b; Smith et al., 2023; Homma et al., 2024; Tan et al., 2024). Furthermore, current and near-future surveys such as the DECam Local Volume Exploration survey (DELVE; Drlica-Wagner et al., 2021, 2022), the Hyper Suprime-Cam Strategic Survey Program (HSC-SSP; Aihara et al., 2018), the Ultraviolet Near Infrared Optical Northern Survey (UNIONS; Ibata et al., 2017), and the Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) are expected significantly increase the known population of MW satellites in the coming years (e.g., Hargis et al., 2014; Jethwa et al., 2018; Newton et al., 2018; Nadler et al., 2020; Manwadkar and Kravtsov, 2022; Nadler et al., 2024a). This growing population of UFDs will improve our understanding of dark matter and allow for more sensitive searches for potential deviation from CDM.